OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4. Answer slope = 3; y intercept at (0, -1) slope = -1; y intercept at (0, -) slope = -3; y intercept at (0, 1) slope = 1; y intercept at (0, -)

OpenStudy (smurfy14):

-3y=-3x+4 y=x-4/3 slope =1

OpenStudy (noseboy908):